OpenStudy (anonymous):

Find the average rate of change of the function over the given interval or intervals. G(t)= 2 + cos t. Interval [0, pi]

OpenStudy (sriram):

2/pi

OpenStudy (anonymous):

that is correct. need the steps?

OpenStudy (sriram):

find value at 0 and then at pi then divide their difference by length of interval

OpenStudy (anonymous):

oh whoops maybe it isn't lets check. average rate is \[\frac{G(\pi)-G(0)}{\pi-0}\] \[=\frac{2+\cos(\pi)-(2+\cos(0))}{\pi}\] \[=\frac{2-1-(2+1)}{\pi}\] \[=\frac{-2}{\pi}\]

OpenStudy (anonymous):

Thank you so much! I actually just solved it too and got -2/pi. Was just making sure it was right.